A person can just survive a full-body collision (either to the front, back, or side) which results in a deceleration that is about 100 g. (One g is 9.8 m/s/s). At greater deceleration fatal brain damage will likely occur. If a 66.6 kg man falls off a cliff of height 35.8 m but manages to land flat on his back in soft snow, undergoing a constant deceleration of this magnitude, how deep would he be buried in the snow?
1
Expert's answer
2018-09-24T12:32:09-0400
v=√2gh
t=mv/F
F=100g
s=vt=(mv^2)/100g=mh/50=(66.6·35.8)/50=47.68 metres